Graph each system of equations. Determine whether the system has no solutions, one solutions, or infinitely many solutions. If the system has one solutions, name it.
3x - y = -2
3y - y = 0

Answer by Alan3354(30958)

3x - y = -2
3y - y = 0 --> 2y = 0
y = 0
-----
3x = -2
x = -2/3